INAF Science

Fixed from yesterday, see IT#8212

01:08
  • Ready to start OSCO
  • Optimize gain failed the first time- drove the mirror too hard.
housekeeper.L |INF| 2709723|2020-10-14 01:22:40.851602| SwitchSafeSkipFrameCnt > 1
housekeeper.L |INF| 2709731|2020-10-14 01:22:41.938999| SwitchSafeSkipFrameCnt > 232
housekeeper.L |INF| 2709739|2020-10-14 01:22:43.037958| SwitchSafeSkipFrameCnt > 597
housekeeper.L |INF| 2709747|2020-10-14 01:22:44.132949| SwitchSafeSkipFrameCnt > 607
housekeeper.L |INF| 2709775|2020-10-14 01:22:45.285118| SwitchSafeSkipFrameCnt > 765
housekeeper.L |INF| 2709787|2020-10-14 01:22:46.379596| SwitchSafeSkipFrameCnt > 797
housekeeper.L |INF| 2709803|2020-10-14 01:22:48.550044| SwitchSafeSkipFrameCnt > 849
housekeeper.L |INF| 2709811|2020-10-14 01:22:49.631534| SwitchSafeSkipFrameCnt > 1022
housekeeper.L |INF| 2709819|2020-10-14 01:22:50.731353| SwitchSafeSkipFrameCnt > 1779
housekeeper.L |INF| 2709827|2020-10-14 01:22:51.816249| SwitchSafeSkipFrameCnt > 2055
housekeeper.L |INF| 2709835|2020-10-14 01:22:52.914910| SwitchSafeSkipFrameCnt > 3797
housekeeper.L |INF| 2709843|2020-10-14 01:22:54.003263| SwitchSafeSkipFrameCnt > 5351
housekeeper.L |INF| 2709871|2020-10-14 01:22:55.202642| SwitchSafeSkipFrameCnt > 5378
housekeeper.L |INF| 2709967|2020-10-14 01:23:05.084363| SwitchSafeSkipFrameCnt > 5829
housekeeper.L |INF| 2709979|2020-10-14 01:23:06.173202| SwitchSafeSkipFrameCnt > 7549
housekeeper.L |INF| 2709987|2020-10-14 01:23:07.254018| SwitchSafeSkipFrameCnt > 9381
housekeeper.L |INF| 2709995|2020-10-14 01:23:08.346511| SwitchSafeSkipFrameCnt > 11218
housekeeper.L |INF| 2710003|2020-10-14 01:23:09.440962| SwitchSafeSkipFrameCnt > 11877
housekeeper.L |INF| 2710091|2020-10-14 01:23:18.206588| SwitchSafeSkipFrameCnt > 12751
housekeeper.L |INF| 2710099|2020-10-14 01:23:19.298911| SwitchSafeSkipFrameCnt > 13871

  • I did it in intervention mode the second time and it was fine...
01:39
  • Let's try this again.
  • Wed Oct 14 01:43:07.779 2020 aos.AcquireRef{FLAO}.failed - AcquireRef{FLAO}[222] - Illegal command for state AOArbitrator.Ready
  • Loop didn't open properly after the last preset.
  • Now we can go.
  • g_tt_0.30_ho1_0.35_ho2_0.30.fits
04:15
  • That went so smoothly that I almost dozed off.
  • We're doing some spectroscopy thingy for a minute.
  • I'm concerned that the LN gains are too low for good correction, so I manually optimized gain.
  • Tip/tilt was very sensitive, so during the very good seeing, I optimized tt to be only 0.15.
  • I suspect that doing an automatic optimize gain will cause the loop to open.
04:20
  • New ESM preset
  • Doing the spectroscopy with ESM. Same gains as before.
  • g_tt_0.15_ho1_0.35_ho2_0.30.fits
  • Tracking Numbers
  • 20201014_043342, 20201014_043440, 20201014_043553, 20201014_043652
-- GregoryTaylor - 14 Oct 2020
Topic revision: r3 - 14 Oct 2020, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback